Write an equation of the line with the following characteristics.
Through (-4,-9),parallel to 4x+15y=-34

Answer by tommyt3rd(5050)

m=-A/B=-4/15
y-(-9)=-4/15 (x-(-4))
y+9 = -4/15 (x+4)